OpenStudy (ray1998):

Please help? Find the sum of the following arithmetic series: 7+12+17+22+...+52 The sum of the series is ___. Do I just add? Thank you in advance! :)

rishavraj (rishavraj):

hmmmm go on u just need to find the no of terms ...and then u can addd :)

OpenStudy (anonymous):

What's the difference between each consecutive number in the sum?

OpenStudy (ray1998):

5

OpenStudy (anonymous):

Right. And what number does the sum stop at?

rishavraj (rishavraj):

so a_1 = 7 and d = 5 and a_n = 52 ....

rishavraj (rishavraj):

and also \[a_n = a_1 + (n - 1)d\] plug the values u got and get n i.e the no of terms...and find the sum of the series

OpenStudy (ray1998):

Yes, I understand that @rishavraj and what do you mean @DurableToaster , the sum of all the numbers?

rishavraj (rishavraj):

so now u can proceeed ??@Ray1998

OpenStudy (ray1998):

7=52 +(5-1)5 is this correct?

rishavraj (rishavraj):

@Ray1998 ???

OpenStudy (ray1998):

Are the numbers I plugged in to the formula correct?

rishavraj (rishavraj):

nope ...see a_1 = 7 , d = 5, a_n = 52 so 52 = 7 + (n - 1)5

OpenStudy (ray1998):

Okay, gotcha. n = 1 + 5 square root 45 n = 3.14112736

OpenStudy (ray1998):

Sorry, my answer wouldn't send in :/

rishavraj (rishavraj):

nope thts wrong .... -_- 52 = 7 + 5n - 5 52 = 2 + 5n 50 = 5n n = 10 so now apply the summation formula

OpenStudy (ray1998):

Sorry, I'm really bad with math. 52 = 7 + (10 - 1)5

rishavraj (rishavraj):

noope.....u need to use now \[S_n = \frac{ n }{ 2 }(a_1 + a_n)\]

rishavraj (rishavraj):

u got all the values now ......solve it @Ray1998

OpenStudy (ray1998):

s = 295/n?

rishavraj (rishavraj):

hmmmm why is tht ...??? :/ \[S_{10} = \frac{ 10 }{ 2 }(7 + 52) \] \[S_{10} = 295 \]

rishavraj (rishavraj):

http://prntscr.com/ajp81h tht part makes no sense